# Connectedness of the Gromov boundary of fine curve graphs

### v2

@@ 3 Fine curve graph
The finite set $P$ can be empty. In that case, it yields an isometric embedding of $\operatorname{\mathcal{NC}}(S-P)\hookrightarrow\operatorname{\mathcal{NC}}^{\dagger}(S)$. This can be realised by choosing a hyperbolic metric on $S$ and each isotopy class of simple closed curves is sent to its geodesic representative.
Now we will fix an arbitrary base point $o\in\operatorname{\mathcal{NC}}^{\dagger}(S)$. We write $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ the subgraph of $\operatorname{\mathcal{NC}}^{\dagger}(S)$ where all the vertices are transverse to $o$. It is a priori equipped with the subgraph distance but not the induced distance.
@@
Let $\alpha\in\operatorname{\mathcal{NC}}^{\dagger}(S)$. Let $F\subset\operatorname{\mathcal{NC}}^{\dagger}(S)$ be a finite collection of pairwise transverse curves. Suppose that for every $\gamma\in F$, $\alpha\cap\gamma\neq\emptyset$. Then there exists a curve $\alpha^{\prime}\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ isotopic to $\alpha$ that is disjoint from $\alpha$ and is transverse to all $\gamma\in F$.
@@
Let $A\subset S$ be a small enough annular neighbourhood of the curve $\gamma\in F$. Since $\alpha$ is a proper embedding of $S^{1}$, there are only finitely many connected components in $\alpha\cap A$. Hence it is possible to cut $A$ into finitely many pieces $A_{1},\ \dots,\ A_{n}$ so that for each $1\leq i\leq n$, the piece $A_{i}$ contains exactly one connected component of $\alpha\cap A$. Note that each $A_{i}$ is diffeomorphic to a $[-1,1]^{2}\subset\mathbb{R}^{2}$ with the image of $o\cap A_{i}$ being $[-1,1]\times\{0\}$. It is an elementary exercise to show that under this diffeomorphism, $\alpha$ is isotopic to an $\alpha^{\prime}$ that is transverse to $\gamma$ and disjoint from $\alpha$ inside of $[-1,1]^{2}$. Note that if $A$ is small enough, then $\alpha^{\prime}$ can be taken so that whenever $\alpha$ is transverse to some other finitely many curves, $\alpha^{\prime}$ remains transverse to them. Hence by induction, we can prove the desired result. ∎
We notice that the perturbation in Lemma 3.7 can be done within an arbitrarily small annular neighbourhood. All the isotopic perturbation in the sequel will be of the same flavour, although the situation might be slightly different.
Proposition 3.8.
The inclusion map $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\hookrightarrow\operatorname{\mathcal{NC}}^{\dagger}(S)$ is a $(1,2)$-quasi-isometry, hence $\partial\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\simeq\partial\operatorname{\mathcal{NC}}^{\dagger}(S)$.
@@
For any $\alpha\in\operatorname{\mathcal{NC}}(S)$, if it is not transverse to $o$, then by Lemma 3.7, there is another $\alpha^{\prime}\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ that is disjoint from $\alpha$ and is transverse to $o$. Hence $d^{\dagger}(\alpha,\alpha^{\prime})=d^{\dagger}\big{(}\alpha,\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\big{)}=1$.
For any two vertices $\alpha,\beta\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$, if they are not transverse, then by applying Lemma 3.7, we can find a $\beta^{\prime}$ with $d^{\dagger}(\beta,\beta^{\prime})=1$ that is transverse to both $\alpha$ and $o$. Now we can find a geodesic $\alpha=\nu_{0},\ \nu_{1},\ \dots,\ \nu_{n}=\beta^{\prime}$ in $\operatorname{\mathcal{NC}}^{\dagger}(S)$ connecting $\alpha$ to $\beta^{\prime}$ with the property that there is a finite subset $P\subset S$ disjoint from $o$ such that $\nu_{i}\cap P=\emptyset$ for all $i$ and $[\alpha]_{S-P}=[\nu_{0}]_{S-P},\ [\nu_{1}]_{S-P},\ \dots,\ [\nu_{n}]_{S-P}=[\beta]_{S-P}$ is also a geodesic in $\operatorname{\mathcal{NC}}(S-P)$. Without loss of generality, we may assume that $\nu_{i}$’s and $o$ are in their minimal positions in $S-P$, by taking a hyperbolic metric making $o$ a geodesic in $S-P$ as per Proposition 3.4, which in particular implies that $\nu_{i}$’s are transverse to $o$ and are contained in $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$. Hence the subgraph distance between $\alpha,\beta^{\prime}\in\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)$ is exactly their distance induced from $\operatorname{\mathcal{NC}}^{\dagger}(S)$.
Now combine the results of two paragraphs above, we can conclude that the inclusion map $\operatorname{\mathcal{NC}^{\dagger}_{o,\pitchfork}}(S)\hookrightarrow\operatorname{\mathcal{NC}}^{\dagger}(S)$ is a $(1,2)$-quasi-isometry and thus by Proposition 2.3, their Gromov boundaries are homeomorphic. ∎

### v4

@@ 3 Fine curve graph and its subgraphs
Using the $d_{\mathcal{C}^{s}(S-P)}$-distance estimation for curve in minimal position instead of $d_{\mathcal{NC}(S-P)}$, similar arguments in Lemma 3.12 also apply to non-separating curves, i.e. for any curve $\alpha\in\mathcal{C}^{\dagger}(S)$ transverse to $o$ with $d^{\dagger}(o,\alpha)=r$, we can find a geodesic path of length $r$ in $\mathcal{C}^{\dagger}(S)$ connecting $o$ to $\alpha$ such that all vertices on the path are curves transverse to $o$.
4 Bounded geodesic image theorem
4.A. Subsurface projection.
Let $S$ be an orientable, connected closed surface of genus $g\geq 2$. Let $Y$ be a compact, proper, non-sporadic subsurface of $S$. Assume that $\partial Y$ is a finite collection of smooth curves in $\mathcal{C}^{\dagger}(S)$.
Let $\alpha\in\mathcal{C}^{\dagger}(S)$ be a simple closed curve on $S$, and let $Y$ be a subsurface as defined above. We say that $\alpha$ intersects $Y$ essentially if at least one connected component of $\alpha\cap Y$ is an essential arc or curve in $Y$, i.e., not isotopically trivial in $Y$. We remark that a general curve $\alpha$ may not intersect $\partial Y$ transversely and this notion of essential intersection may not be invariant under isotopy.
If $\alpha$ intersects $Y$ essentially, then fix an essential arc $\tau\subset\alpha\cap Y$ and denote by $c_{\pm}\subset\partial Y$ the two boundary components where the endpoints of $\tau$ are located. Note that $c_{+}$ and $c_{-}$ might coincide. Let $\mathcal{N}$ be a regular neighbourhood of $\tau\cup c_{\pm}$ in $Y$ with a non-trivial smooth boundary, i.e. a boundary that is an essential curve in $Y$.
Figure 1: A possible subsurface projection $\pi_{Y}(\alpha)$ of curve $\alpha$ on subsurface $Y$.
Definition 4.1 (Subsurface projection).
Suppose that $S$ and $Y$ are given as above. Let $\mathcal{C}^{\dagger}(Y)$ be the fine curve graph for $Y$ and let $\mathcal{P}\left(\mathcal{C}^{\dagger}(Y)\right)$ be its power set. We define a map $\pi_{Y}\colon\mathcal{C}^{\dagger}(S)\to\mathcal{P}\left(\mathcal{C}^{\dagger}(Y)\right)$ in the following way: for each $\alpha\in\mathcal{C}^{\dagger}(S)$, the image $\pi_{Y}(\alpha)$ is defined as
  • $\{\alpha\}$ if $\alpha\subset Y$;
  • $\emptyset$ if $\alpha$ does not intersect $Y$ essentially;
  • the union of all $\partial\mathcal{N}\cap Y$ if $a\not\subset Y$ but $\alpha$ intersects $Y$ essentially, where $\mathcal{N}$ is a regular neighbourhood as above.
The map $\pi_{Y}$ is called the subsurface projection of $\alpha$ on $Y$.
Nevertheless, this map is coarsely well-defined:
Proposition 4.2.
With the simplicial metric $\big{(}\mathcal{C}^{\dagger}(Y),d^{\dagger}_{Y}\big{)}$, for any curve $\alpha\in\mathcal{C}^{\dagger}(S)$, the diameter of the set $\pi_{Y}(\alpha)$ is bounded in $\mathcal{C}^{\dagger}(Y)$, i.e.,
$\operatorname{diam}_{d^{\dagger}_{Y}}(\pi_{Y}(\alpha))\leq 11.$
Proof.
It suffices to prove the case where $\alpha$ intersects $Y$ essentially but is not contained in $Y$, as other situations are trivial. Let $\tau_{1}$ and $\tau_{2}$ be two essential arcs in $\alpha\cap Y$. We note that $\tau_{1}$ and $\tau_{2}$ are either identical or disjoint. Denote by $c^{1}_{\pm}$ and $c^{2}_{\pm}$ the boundary components of $Y$ where the starting and ending points of $\tau_{1}$ and $\tau_{2}$ are located, respectively. For $i=1,2$, let $\gamma_{i}$ be a boundary component of a regular neighbourhood $\mathcal{N}_{i}$ of $\tau_{i}\cup c^{i}_{\pm}$.
We can choose two smaller regular neighbourhoods $\widehat{\mathcal{N}_{i}}$ of $\tau_{i}\cup c^{i}_{\pm}$ for $i=1,2$, with smooth boundary components within $Y$, such that $\widehat{\mathcal{N}_{i}}\subset\mathcal{N}_{i}$ and $\partial\widehat{\mathcal{N}_{i}}$ and $\partial\widehat{\mathcal{N}_{j}}$ are in minimal position for $i\neq j$. Let $\widehat{\gamma_{i}}$ denote a boundary component of $\partial\widehat{\mathcal{N}_{i}}$. Due to disjointness, we have $d_{Y}^{\dagger}(\widehat{\gamma_{i}},\gamma_{i})=1$ for $i=1,2$.
Observe that $|\widehat{\gamma_{1}}\cap\widehat{\gamma_{2}}|\leq 4$. Since $\widehat{\gamma_{1}}$ and $\widehat{\gamma_{2}}$ are in minimal position within $Y$, we apply Proposition 3.8 to obtain
$d_{Y}^{\dagger}(\gamma_{1},\gamma_{2})\leq d_{Y}^{\dagger}(\gamma_{1},\widehat{\gamma_{1}})+d_{Y}^{\dagger}(\widehat{\gamma_{1}},\widehat{\gamma_{2}})+d_{Y}^{\dagger}(\gamma_{2},\widehat{\gamma_{2}})\leq 11.$
This completes the proof. ∎
In addition to Proposition 4.2, we have several other distance estimates for curves that intersect the subsurface essentially and their corresponding subsurface projections.
Proposition 4.3.
Let $S$, $Y$, $d^{\dagger}_{Y}$, and $\pi_{Y}$ be as defined above. For any two curves $\alpha,\beta\in\mathcal{C}^{\dagger}(S)$ that intersect the subsurface $Y$ essentially with $|\alpha\cap\beta|=p\geq 0$, and for any $\gamma_{\alpha}\in\pi_{Y}(\alpha)$ and $\gamma_{\beta}\in\pi_{Y}(\beta)$, the following inequality holds:
$d_{Y}^{\dagger}\big{(}\gamma_{\alpha},\gamma_{\beta}\big{)}\leq 8p+11.$
Proof.
By choosing the regular neighbourhoods in Definition 4.1 sufficiently small, we can find curves $\widehat{\gamma_{\alpha}}\in\pi_{Y}(\alpha)$ and $\widehat{\gamma_{\beta}}\in\pi_{Y}(\beta)$ such that they are transverse within $Y$ and satisfy $|\widehat{\gamma_{\alpha}}\cap\widehat{\gamma_{\beta}}|\leq 4+4p$. Additionally, we ensure that $\gamma_{\alpha}\cap\widehat{\gamma_{\alpha}}=\emptyset$ and $\gamma_{\beta}\cap\widehat{\gamma_{\beta}}=\emptyset$, which implies $d_{Y}^{\dagger}(\gamma_{\alpha},\widehat{\gamma_{\alpha}})=d_{Y}^{\dagger}(\gamma_{\beta},\widehat{\gamma_{\beta}})=1$. By applying Proposition 3.8, we obtain $d_{Y}^{\dagger}\big{(}\widehat{\gamma_{\alpha}},\widehat{\gamma_{\beta}}\big{)}\leq 8p+9$. Then, using the triangle inequality, we conclude that $d_{Y}^{\dagger}\big{(}\gamma_{\alpha},\gamma_{\beta}\big{)}\leq 8p+11$. ∎
Using the triangle inequality and applying Proposition 4.3, we can further deduce the following corollary:
Corollary 4.4.
Let $P=\left\{\alpha_{0},\alpha_{1},\ldots,\alpha_{n}\right\}$ be a sequence in $\mathcal{C}^{\dagger}(S)$ and $p>0$ be a positive integer. Suppose that every $\alpha_{i}$ intersects $Y$ essentially and that $|\alpha_{i}\cap\alpha_{i+1}|\leq p$, then $d^{\dagger}_{Y}\left(\pi_{Y}(\alpha_{0}),\pi_{Y}(\alpha_{n})\right)\leq(8p+11)n$.
4.B. Surgery.
To prove Theorem 1.3, we will employ the auxiliary process of subsurface surgery, similar to the classical cases. The idea is to adapt the proof from [webb2013shortproofboundedgeodesic] to the case of the fine curve graph.
First, let us introduce the following notion:
Definition 4.5 (Sensible collection).
Let $S$ be a hyperbolic surface, possibly with punctures. We say that a collection of essential multicurves $\left(\gamma_{i}\right)\subset\mathcal{C}^{\dagger}(S)$ is sensible if the curves are essential and have no triple points, i.e. for distinct indices $i,j,k$, we have $\gamma_{i}\cap\gamma_{j}\cap\gamma_{k}=\emptyset$.
Throughout this section, we will assume that $\alpha,\beta\in\mathcal{C}^{\dagger}(S)$ are two curves in minimal position that fill $S$, i.e. cut $S$ into a disjoint union of discs.
Recall that when we orient $\gamma$ and $\beta$, each point in $\gamma\cap\beta$ is assigned an intersection sign of ±$\pm 1$. We say that a pair of such points have opposite signs if their signs differ, and the same sign otherwise. Note that this notion does not depend on the choice of orientation.
Definition 4.6 (Loop).
Let $S$ be a surface and $\alpha,\beta$ be as above. We say that $\gamma$ is an $(\alpha,\beta)$-loop if for each subarc $b\subset\beta-\alpha$, we have $|\gamma\cap b|\leq 2$, with equality only if $\gamma\cap b$ have opposite sign.
In the following paragraphs, we will assume only that $\alpha$ is a multicurve, while $\beta$ and $\gamma$ are essential curves on $S$. Given a sensible triplet $\{\gamma,\alpha,\beta\}$, we aim to construct a new curve $\chi(\gamma)$ that forms an $(\alpha,\beta)$-loop if $\gamma$ does not already do so.
Suppose now that $\{\gamma,\alpha,\beta\}$ is a sensible triplet and that $\gamma$ is not an $(\alpha,\beta)$-loop. Let $c$ be a minimal connected subarc of $\gamma$, with respect to inclusion, such that there exists an arc $b\subset\beta-\alpha$ satisfying one of the following conditions:
  • $c\cap b$ consists of a pair of points with the same sign, or
  • $|c\cap b|\geq 3$.
Since $c$ is minimal and connected, this implies that $c$ has endpoints on $b$, and that $b$ is the unique arc satisfying the above properties, with the additional condition that $|c\cap b|\leq 3$.
Now the intersection between $b$ and $c$ falls into the following three cases, with the same surgery as in [webb2013shortproofboundedgeodesic] recorded here below. Let $\mathcal{N}$ be a small enough regular neighbourhood of $\beta$ and $\mathcal{R}\subset\mathcal{N}-\alpha$ a rectangle containing $b$.
  1. (K1)
    $|c\cap b|=2$ with same sign. Then let $\{p_{1},p_{2}\}=c\cap\partial\mathcal{R}$. Connect $p_{1}$ to $p_{2}$ (or vice versa) by an arc $a\subset\mathcal{R}$ that intersects $b$ only once and intersects $c$ only at the endpoints of $a$. We let $\chi(\gamma)$ be the simple closed curve $a\cup(c-\mathcal{R})$. See Figure 2
  2. (K2)
    $|c\cap b|=3$ with alternating intersection signs with respect to some order on $b$. Let $p_{1},p_{2},p_{3}$ be the three points $c\cap b$ in some order along $b$. Let $c_{1},c_{2}\subset c$ be arcs such that $c_{1}\cup c_{2}=c$, $\partial c_{1}=\left\{p_{1},p_{2}\right\}$ and $\partial c_{2}=\left\{p_{2},p_{3}\right\}$. Connect $c_{1}\cap\partial R$ to $c_{2}\cap\partial\mathcal{R}$ by two disjoint arcs $a_{1},a_{2}\subset\mathcal{R}$ so that $a_{1}$ intersects $c_{1},c_{2}$ only at its endpoints and intersects $b$ once, and similarly $a_{2}$. We let $\chi(\gamma)=a_{1}\cup\left(c_{1}-\mathcal{R}\right)\cup a_{2}\cup\left(c_{2}-\mathcal{R}\right)$.
  3. (K3)
    $|c\cap b|=3$ with non-alternating intersection signs. Let $p_{1},p_{2},p_{3}$ be the three points $c\cap b$ in some order along $b$. Without loss of generality, assume that $p_{1}$ and $p_{2}$ have the same sign. Then connect $p_{1}$ to $p_{2}$ with an arc $a$ similarly to (K1) so that the sign at the unique intersection point between $a$ and $b$ is different from the intersection sign of $p_{3}$.
We denote by $\chi(\gamma)$ this curve newly constructed from $\gamma$ according to the aforementioned procedures.
Figure 2: Surgery for (K1) (left) and (K3) (right).
Now let $S$ be again an oriented closed surface with genus $g(S)\geq 2$. Let $\beta\in\mathcal{C}^{\dagger}(S)$ be a curve and $\alpha$ be a multicurve that is transverse to $\beta$. Suppose that $P\subset S$ is a finite collection of points on the surface such that $\alpha,\beta$ are in minimal position on $S-P$. Moreover, we assume that $p\in P$ if and only if $p$ is contained in a bigon between $\alpha$ and $\beta$ and that each bigon contains only one point in $P$.
@@ 4 Bounded geodesic image theorem
Let $S$, $\alpha$, $\beta$ and $P$ be as above. Suppose that $\gamma\in\mathcal{C}^{\dagger}(S)$ is a curve such that the triplet $(\gamma,\alpha,\beta)$ is sensible on $S-P$. Then we can find $\chi(\gamma)\in\mathcal{C}^{\dagger}(S-P)$ is an $(\alpha,\beta)$-loop and is surviving, i.e. $\chi(\gamma)\in\mathcal{C}^{\dagger}(S)$.
@@
The same proof in [webb2013shortproofboundedgeodesic, Lemma 2.2] also applies here, resulting that $\chi(\gamma)$ is an $(\alpha,\beta)$-loop and remains essential on $S-P$. To show that $\chi(\gamma)$ is surviving, it suffices to show that $\chi(\gamma)$ can be taken so that they do not only bound a disk in $S$ containing some points in $P$.
For the cases (K1) and (K3), by proof of contradiction, we assume that $\chi(\gamma)$ is not surviving. Then $\chi(\gamma)$ encloses a punctured disk $\Delta$. There are two cases: $\gamma_{1}\subset\Delta$ or $\gamma_{2}\subset\Delta$. In either case, the extension of $\gamma_{i}$ would intersect with $\partial\Delta$. This contradicts the fact that $\gamma$ is a simple closed curve.
For the case (K2), assume also ab absurdo that $\chi(\gamma)$ is not surviving. Observe that $\chi(\gamma)$ has to be has to be separating, since a non-separating curve cannot cut an oriented surface into a disk. This means that the two subsurfaces $B_{\pm}$ bounded $c_{1},c_{2}$ and $b$ are two disks on $S$, see Figure 3. Moreover, by non-surviving condition, it is implied that $B_{-}$ and $B_{+}$ both contains points in $P$. By our assumption on $P$, this further indicates that both $B_{-}$ and $B_{+}$ are contained in a bigon between $\beta$ and $\alpha$. As a result, it forces a component of $\alpha$ to bound a two bigons and become non-essential on $S$, which yields a contradiction. ∎
Figure 3: Surgery for (K2) and its proof of Lemma 4.7.
4.C. Proof of bounded geodesic image theorem.
Since the surgery is now well-defined in the context of the fine curve graph, to prove the bounded geodesic image theorem for the fine curve graph, it suffices to proceed through the lemmata below, following the arguments in [webb2013shortproofboundedgeodesic].
We note that the surgery for $\gamma$ is performed exactly as in [webb2013shortproofboundedgeodesic]. Hence we may also conclude the following generalised result from [hensel2013slimunicornsuniformhyperbolicity, Proposition 3.1.7]:
Lemma 4.8 (cf. Lemma 2.3, [webb2013shortproofboundedgeodesic]).
Let $S$, $P$, $\alpha$ and $\beta$ be as above. Let $\gamma_{1},\gamma_{2}\in\mathcal{C}^{\dagger}(S)$ be two curves such that $\gamma_{1},\gamma_{2},\alpha,\beta$ are sensible on $S-P$. Assume in addition that $\gamma_{1}$ is disjoint from $\gamma_{2}$. Then for any $\chi(\gamma)$ and $\chi(\beta)$ constructed above, we have then $|\chi(\gamma_{1})\cap\chi(\gamma_{2})|\leq 4$.
The above result will allow us to deduce the following distance estimation between the subsurface projection of $(\alpha,\beta)$-loops obtained from two disjoint curves from Proposition 3.8:
Lemma 4.9.
Let $S$, $P$, $\alpha$, $\beta$, $\gamma_{1}$ and $\gamma_{2}$ be as above. Then $d^{\dagger}\big{(}\chi(\gamma_{1}),\chi(\gamma_{2})\big{)}\leq 9$.
With results provided above, following the arguments of [webb2013shortproofboundedgeodesic, Lemma 2.4], we can obtain the following result:
Lemma 4.10.
Let $\alpha^{\prime}$ be a component of a multicurve $\alpha$ on $S$ and let $\beta$ be a curve on $S$. Suppose $\alpha^{\prime},\beta$ are sensible and fill $S$. Then there exists a $(9,0)$-quasigeodesic $\alpha^{\prime}=\gamma_{0},\gamma_{1},\ldots,\gamma_{n}=$ $\beta$ with $\gamma_{i}$ is a $(\alpha,\beta)$-loop for every $0<i<n$.
Moreover, we have the following result:
Lemma 4.11.
Let $Y$ be a non-annular subsurface of $S$ and $\beta\in\mathcal{C}^{\dagger}(S)$. Suppose in addition that a component $a\subset\partial Y$ and $\beta$ fill $S$. Let $\gamma\in\mathcal{C}^{\dagger}(S)$ be a $(\partial Y,\beta)$-loop that intersects $\partial Y$. Then for any $\widehat{\gamma}\in\pi_{Y}(\gamma)$ and $\widehat{\beta}\in\pi_{Y}(\beta)$, we have $d_{Y}(\widehat{\gamma},\widehat{\beta})\leq 24$.
@@
Since $Y$ is non-annular subsurface, any pair of arcs in the respective subsurface projection will intersect at most twice by Definition 4.6, so one can consider a closed regular neighbourhood of the arcs to prove that there are some $\gamma^{\prime}\in\pi_{Y}(\gamma)$ and $\beta^{\prime}\in\pi_{Y}(\beta)$ satisfying $d_{Y}(\gamma^{\prime},\beta^{\prime})\leq 2$. Applying triangle inequality and Proposition 4.2, we can conclude the result. ∎
With the lemmata above, we can now prove Theorem 1.3.
Proof of Theorem 1.3.
Since we have finished all the preparation using surgery for the proof of bounded geodesic image theorem. Note that the fine curve graph is also Gromov hyperbolic, the argument from [webb2013shortproofboundedgeodesic] also applies here to conclude Theorem 1.3. ∎
4.D. A surviving version.
Assume that $S$ is a surface as given above and $P\subset S$ is a finite collection of point on $S$. Let $Y\subset S-P$ be a proper subsurface that is not sporadic. We consider the subsurface projection $\widetilde{\pi_{Y}}\colon\mathcal{C}^{s}(S-P)\to\mathcal{P}(\mathcal{C}(Y))$ given as in the classical cases, i.e. the image of an (isotopy class of a) surviving curve is the isotopy class of curves defined as in §4.A.
We observe that the only arguments in the proof of the bounded geodesic image theorem that depend on the topological type of the curves are the surgeries performed in Proposition 3.8 and in §4.B. As shown in Proposition 3.8 and Lemma 4.7, the new curves obtained from the surgeries remain surviving on $S-P$. Recall that the surviving curve graph $\mathcal{C}^{s}(S-P)$ is also Gromov hyperbolic (see [rasmussen2020uniform] and [bowden2022quasi, theorem 3.3]). Following the approach in [webb2013shortproofboundedgeodesic], we conclude:
Theorem 4.12.
Let $S$ be a closed surface and $P\subset S$ be a finite collection of point. Then there exists an $M\geq 0$ such that for any non-sporadic proper subsurface $Y\subset S-P$ and any geodesic $(g_{i})\subset\mathcal{C}^{s}(S-P)$ where each $g_{i}$ intersects $Y$ essentially, we have $\operatorname{diam}_{d_{Y}}\big{(}\widehat{\pi_{Y}}(g_{i})\big{)}\leq M$.
Remark 4.13.
As both $\pi_{Y}$ and $\widehat{\pi_{Y}}$ are coarsely well-defined for a curve $\alpha$ intersecting $Y$ essentially, we will simply write $\pi_{Y}(\alpha)$ and $\widehat{\pi_{Y}}(\alpha)$ for an (arbitrary) element in the projection.
4.E. Application: non-compactness.
As an interesting application of bounded geodesic image theorem, we can show that the Gromov boundary and the Gromov bordification of the fine curve graph are not compact.
Proof of Theorem 1.4.
Let $o\in\operatorname{\mathcal{NC}}^{\dagger}(S)$ and let $\xi\in\partial\mathcal{C}^{\dagger}(S)$ be a point on the Gromov boundary. Let $Y=S-o$ be the complementary component and $a\in\mathcal{C}^{\dagger}(S)$ that intersects $Y$ essentially and $\langle\xi,a\rangle_{o}\gg 1$. Take $\varphi\in\mathrm{Homeo}(S)$ be a homeomorphism on $S$ such that $\varphi|_{Y}$ is pseudo-Anosov on $Y$.
We claim that the sequence $\big{(}\varphi^{n}(\xi)\big{)}_{n\geq 1}$ does not have an accumulation point on $\partial\mathcal{C}^{\dagger}(S)$. Since $\partial\mathcal{C}^{\dagger}(S)$ is metrisable, suppose for contradiction that there is a convergent subsequence $\big{(}\varphi^{n_{k}}(\xi)\big{)}_{k\geq 1}$. By (2.3), this implies that for any $R>0$, there exists a sufficiently large $k_{0}>0$ so that for an arbitrary $k\gg k_{0}$, we have $\langle\varphi^{n_{k_{0}}}(a),\varphi^{n_{k}}(a)\rangle_{o}>R$. We may assume that $\varphi^{n_{k_{0}}}(a),\varphi^{n_{k}}(a)$ are transverse to $o$ (as such a change causes an additive error at most $2$ in the Gromov product). Let $P\subset S$ such that $\varphi^{n_{k_{0}}}(a)$, $\varphi^{n_{k}}(a)$ and $o$ are pairwise in minimal position. Since $k\gg k_{0}$ and $\varphi$ is pseudo-Anosov on $Y$, we have then
$d_{Y}\big{(}\pi_{Y}\big{(}\varphi^{n_{k_{0}}}(a)\big{)},\big{(}\varphi^{n_{k}}(a)\big{)}\big{)}>M,$
where $M>0$ is the constant from Theorem 4.12. Appealing to Theorem 4.12, this means that any geodesic connecting $[\varphi^{n_{k_{0}}}(a)]_{S-P}$ to $[\varphi^{n_{k}}(a)]_{S-P}$ must contain an isotopy class of curve that does not intersect $Y$ essentially, which has to be $[o]_{S-P}$. So by Lemma 3.7, we can compute
$\langle\varphi^{n_{k_{0}}}(a),\varphi^{n_{k}}(a)\rangle_{o}$ $=\frac{1}{2}\left(d^{\dagger}\big{(}\varphi^{n_{k_{0}}}(a),o\big{)}+d^{\dagger}\big{(}\varphi^{n_{k}}(a),o\big{)}-d^{\dagger}\big{(}\varphi^{n_{k_{0}}}(a),\varphi^{n_{k}}(a)\big{)}\right)$
$=\frac{1}{2}\Big{(}d_{C^{s}(S-P)}\big{(}[\varphi^{n_{k_{0}}}(a)]_{S-P},[o]_{S-P}\big{)}+d^{\dagger}\big{(}[\varphi^{n_{k}}(a)]_{S-P},[o]_{S-P}\big{)}$
$\qquad-d^{\dagger}\big{(}[\varphi^{n_{k_{0}}}(a)]_{S-P},[\varphi^{n_{k}}(a)]_{S-P}\big{)}\Big{)}$
$=0<R,$
contradiction! Hence $\partial\mathcal{C}^{\dagger}(S)$ is not sequentially compact. As it is a metrisable space, it is not compact. ∎
Note that the bordification of the fine curve graph is also completely metrisable (cf. [das2017geometry, Corollary 3.6.14]). So if it is compact, it will also be sequentially compact. But the sequence $\big{(}\varphi^{n}(a)\big{)}_{n\geq 1}$ as in the proof of Theorem 1.4 does not contain a subsequence converging in in $\mathrm{bord}\,\mathcal{C}^{\dagger}(S)$, otherwise the Gromov product between two elements from the subsequence has to converge in $(0,\infty]$, which is not the case. The existence of such a sequence implies again that the bordification is not compact:
Corollary 4.14.
Let $S$ be an orientable connected closed surface with genus $g\geq 2$. Then the Gromov bordification $\mathrm{bord}\,\mathcal{C}^{\dagger}(S)$ is not compact.